OpenStudy (xmoses1):

Which of the following is not equal to 485 L? 0.485 kL 4.85 x 104 mL 48,500 cL 4.85 x 108 µL

OpenStudy (johnweldon1993):

\[\large 485L \cancel{=} 4.85 \times 10^4mL\] \[\large 485L = 4.85 \times 10^5mL\]

OpenStudy (xmoses1):

Thanks!

OpenStudy (johnweldon1993):

No problem!